wolod io 15) Provide the major organic product which results when (CH3)2CHCH2CH2cocI is treated with LIAIH[ OC (CH3)3 3.
Answer #1

= 24 CH) CH-CH, CH - CA hệ 557, mechanism (CH3 -CH=CH-CHE (CH₂) =CH-CH=CH-C-H 4* Al [OCCcm)] - This reagent is highly specifi